OpenStudy (anonymous):

what is derivative of Ux and Uy of U = x^(1/2)*y^(1/2)

OpenStudy (anonymous):

partials?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for d/dx (f(x,y)), just treat any "y's" as constants or d/dy (f(x,y)), just treat any "x's" as constants

OpenStudy (anonymous):

If i divide derivative of Ux and Uy would it be 1?

OpenStudy (anonymous):

eg.: d/dx ( y) = 0 d/dx (y*x) = y d/dx (y^2*x) = y^2 d/dx(y^2*x^2) = y^2 * 2x

OpenStudy (anonymous):

These two are exanmples that might help you better understand what @Algebraic! means.